# The X-ray rise and fall of the Recurrent Symbiotic System T CrB

### v1

@@ Abstract
We present the analysis of publicly available NuSTAR, Suzaku and XMM-Newton observations of the recurrent symbiotic system T CrB covering the 2006.77–2022.66 yr period. The X-ray spectra are analysed by adopting a model that includes a reflection component produced by the presence of the accretion disk. Our best-fit model requires a disk with a radius of 100 AU, effective thickness of 10 AU, averaged column density 1025 cm-2 and orientation of 50 with respect to the line of sight that contributes significantly to the total X-ray flux detected from T CrB, in addition to naturally producing the emission of the 6.4 keV Fe line. The spectral analysis suggests that the temperature of the boundary layer evolved from 14.8 keV in the steady-state phase (before 2016), to 2.8 keV in the 2017.24 epoch, to finally stabilise to about $\sim$8 keV in the subsequent epochs. Variations in the plasma temperature of the boundary layer are attributed to the evolution of the mass accretion rate using theoretical interpretations from an accretion disk model. The presence of emission lines in the high-resolution XMM-Newton RGS spectrum of 2017.24 prevents from adopting a black body emission model to fit the soft X-ray range. Instead, we use plasma emission models that suggest the presence of adiabatically-shocked gas produced by gas velocities of 110–200 km s-1, very likely tracing jet-like ejections similar to what is found in other symbiotic systems. Given the the analysis of X-ray and optical data together, we show that T CrB has a similar evolution as black hole binaries, accreting neutron stars and AGN in the hardness-intensity diagram.
@@
pubyear: 2024pagerange: The X-ray rise and fall of the Recurrent Symbiotic System T CrBReferences
@@ 1 Introduction
As a consequence of these works, we have started a series of papers to address the X-ray properties of the reflection component in highly variable symbiotic stars. We want to test further our ionised disk model to try to push forward our understanding of the accretion and reflection processes in symbiotic systems. Here we present the analysis of publicly available X-ray observations of T CrB. This is a recurrent symbiotic star in which the secondary is a M4III star (Mürset & Schmid, 1999). It has a period is 227.6 d (Kenyon & Garcia, 1986; Fekel et al., 2000) and its distance has been estimated to be $d$=896$\pm$22 pc using Gaia data (Bailer-Jones et al., 2021) [^1] Parallax distance measurements of symbiotic stars are affected by significant uncertainties given their extension, colours and orbital motion as pointed out by Linford et al. (2019), however, with all this caveats, the one reported is the best available estimate of T CrB distance..
@@
T CrB was first detected in X-rays by the Einstein observatory (Cordova et al., 1981) and has been routinely observed by more modern X-ray instruments. The first hard X-ray detection of T CrB during its quiescent state was reported by Tueller et al. (2005), but subsequent analysis of the Swift and Suzaku data suggested the presence of heavily-absorbed, high-temperature plasma, producing hard X-ray emission (Luna et al., 2008; Kennea et al., 2009). This source was then classified as a $\delta$-type symbiotic star, but by early 2014 it started brightening in the optical, entering a "super active" phase (Munari et al., 2016). In fact, the optical light curve exhibited a peak around 2016 April 4–5 (2016.35) and subsequent XMM-Newton observations obtained during 2017 showed that its spectrum developed a bright soft component, which changed T CrB classification into a $\beta/\delta$-type (Luna et al., 2018). Further analysis of 2017 and 2018 XMM-Newton observations in comparison with Suzaku and Swift data presented by (Zhekov & Tomov, 2019) showed that the flux of the soft component was declining while the opposite occurred to the hard component. These authors estimated mass accretion rates ($\dot{M}_{\mathrm{acc}}$) for the different epochs and found that this was higher for the quiescent state than for the active phase (see fig. 6 in Zhekov & Tomov, 2019).
@@
T CrB offers us a unique possibility to study the extreme changes of accretion states experienced by a recurrent symbiotic star. In this paper, we present the analysis of multi-epoch (2006.77–2022.66), multi-instrument (XMM-Newton, Suzaku and NuSTAR) X-ray observations that cover the evolution from the steady-state phase up to the current active phase that, in combination with optical monitoring, can help us assess the role of the accretion process and the reflection physics. The analysis presented here is paving the road for the understanding of such an iconic recurrent symbiotic system on its way to producing a nova-like outburst expected in September of the present year [^2] https://blogs.nasa.gov/Watch_the_Skies/2024/02/27/view-nova-explosion-new-star-in-northern-crown/.
@@ 2.1 XMM-Newton observations
Figure 2: $B$- (top panels) and $V$-magnitude (bottom panels) light curves of T CrB. Data from AAVSO, Schaefer (2010) and DASCH are illustrated with different colours. The data from Schaefer (2010) and DASCH have been displaced by 78.47 yr to fit the AAVSO light curve. The vertical lines show the epoch of the X-ray observations analysed here. The grey-shaded area in the rightmost panels represents the possible next nova-like event predicted by Schaefer (2023).
@@ 2.4 Complementary observations
We also retrieved the optical light curves of T CrB from the American Association of Variable Star Observers (AAVSO) [^7] https://www.aavso.org/ obtained with the optical $B$ and $V$ filters which are contemporaneous with the Observations obtained with the current generation of X-ray satellites. The data correspond to photometric measurements obtained between 2004 Mar 11 and 2024 Apr 25. The resultant light curves are shown in Fig. 2. This figure also shows the photometric measurements obtained from the Digital Access to a Sky Century @ Harvard (DASCH) [^8] http://dasch.rc.fas.harvard.edu/lightcurve.php and values listed in Schaefer (2010). The latter data sets were shifted +78.47 yr to make them comparable with the AAVSO data.
@@ 3 Data analysis and Results
It is interesting to note that the RGS spectrum obtained on 2017.24 shows the presence of emission lines (top right panel of Fig. 1). There seems to be emission from the C v, S xi, N vi, Ar xv and N vii at 35.0 Å, 31.5 Å, 29.1 Å, 27.05 Å  and 24.78 Å, respectively, which have been also detected in other X-ray-emitting symbiotic systems (see Toalá, Botello & Sabin, 2023; Toalá et al., 2023). Given that the RGS instruments are best suited to study the emission in the 0.33–2.5 keV energy range, the 2017.24 RGS spectrum is evidently tracing the contribution from the soft component. The presence of these emission lines is also suggested by the medium-resolution EPIC MOS spectra (not shown here) but with evidently lower spectral quality. This suggests that the soft X-ray emission has a thermal origin, instead of a black body as suggested in previous analysis of the XMM-Newton data (see Luna et al., 2018; Zhekov & Tomov, 2019).
@@
$<F_{V}>$ erg cm-2 s-1 7.4×$\times 10^{-10}$ 1.0×$\times 10^{-9}$ 1.2×$\times 10^{-9}$ 9.9×$\times 10^{-10}$ $10\times 10^{-10}$×$\times 10^{-9}$ 1.0×$\times 10^{-9}$
$<F_{B}>$ erg cm-2 s-1 2.8×$\times 10^{-10}$ 6.1×$\times 10^{-10}$ 7.1×$\times 10^{-10}$ 4.5×$\times 10^{-10}$ 5.7×$\times 10^{-10}$ 5.7×$\times 10^{-10}$ 5.5×$\times 10^{-10}$
@@ 3.1 Spectral analysis
The spectral analysis of the X-ray data of T CrB presented here was performed by means of the X-Ray Spectral Fitting Package (xspec; Arnaud, 1996). Following previous analysis of the X-ray data of T CrB we decided to model the heavily-extinguished hot component by adopting an optically-thin emission model apec, included in xspec [^9] https://heasarc.gsfc.nasa.gov/xanadu/xspec/manual/node134.html. However, the presence of emission lines in the soft X-ray range unveiled by the 2017.24 data prevents us from modelling the soft range (0.3–2.0 keV) with a black body emission model. The soft X-ray emission was hence modelled with a second apec component. The extinction is taken into account by adopting photoelectric absorption using cross-sections of the phabs model, also part of xspec.
@@
Taking advantage of the analysis of the $\beta/\delta$ symbiotic system CH Cyg presented by our group, we decided to use tailor-made reflection models in the spectral analysis of T CrB. In Toalá et al. (2023) we produced 40 disk models using the radiative-transfer code reflex (Paltani & Ricci, 2017) which simulates physical processes of X-ray photons propagating through matter around the central source, adopting Monte Carlo calculations to track individual photons. We use the same models here, which are characterised by a radius $R$, a thickness $h$ and an averaged column density $N_{\mathrm{H,ref}}$. For each disk model, we created 10 spectral energy distributions (SED) adopting angles between 0 and 90 (see illustration in Fig. 3). Finally, we converted these models to an additive one-parameter table (in fits format) associated with all the SEDs using the HEASoft task ftflx2tab [^10] https://heasarc.gsfc.nasa.gov/lheasoft/help/ftflx2tab.html. For simplicity, hereafter we will refer to this model as reflection.
Consequently, we adopted the following model configuration:
@@
We used the 2017.24 epoch to restrict $N_{\mathrm{H1}}$ and used this as a fixed value for the rest of the epochs. Table 2 shows that $N_{\mathrm{H1}}$ resulted in 2.3×$\times 10^{21}$ cm-2, which is very close to that estimated by Galactic H i column density towards T CrB (see section 3.2 of Zhekov & Tomov, 2019). The CF was left as a free parameter in the determination of the 2006.77+2015.81 epoch, which resulted in 0.998$\pm$0.001. This value was taken as a fixed parameter for the rest of the epochs.
Figure 6: Top panels: Evolution with time of the X-ray flux $F_{\mathrm{X,TOT}}$ and the contribution from the different model components (left) and the mass accretion rate $\dot{M}_{\mathrm{acc}}$ (right). Bottom: Evolution of the ratio of the X-ray flux over the optical fluxes from the $V$ and $B$ bands as a function of time (left) and $\dot{M}_{\mathrm{acc}}$. The vertical dotted line shows the time of the maximum optical peak reported on 2016.35.
The temperature of the soft component ($kT_{1}$) is more or less consistent within different epochs, with an averaged value of 0.03 keV (=3.5×$\times 10^{5}$ K), however, its observed flux varies about two orders of magnitudes between epochs, being that of 2017.24 the brightest soft component. The heavily-extinguished apec ($kT_{2}$) and the reflection components jointly fit the unresolved Fe feature and, in the case of the 2006.77+2015.81 and 2020.77 epochs, the energy range above 10 keV detected by NuSTAR (see top right panel of Fig. 4 and top left panel of Fig. 5). The reflection model dominates for $E>$15 keV, leveraging the role of reflection in the hard energy range. The heavily-extinguished thermal component shows a temperature ranged from $kT_{2}$=2.8 keV (=3.2×$\times 10^{7}$ K) to values of $kT_{2}$=14.8 keV (=1.7×$\times 10^{8}$ K) with an hydrogen column density varying from $N_{\mathrm{H2}}$=25×$\times 10^{22}$ cm-2 to about 60×$\times 10^{22}$ cm-2. Fig. 6 (top left panel) illustrates the variation of the total intrinsic flux $F_{\mathrm{X,TOT}}$ and the corresponding contributions from the three components ($F_{\mathrm{X,soft}}$, $F_{\mathrm{X,Hard}}$ and $F_{\mathrm{X,Ref}}$).
@@ 4 Discussion
As mentioned before, using the 2006.77 Suzaku spectrum it was possible to assign a $\delta$-type to T CrB (see table 1 in Merc et al., 2019), but any other latter medium-resolution X-ray spectrum can undoubtedly be used to classify this system as a $\beta/\delta$-type. This strongly suggest that the $\alpha$, $\beta$, $\delta$, and $\beta/\delta$ classification scheme, initially proposed by Mürset et al. (1997) and expanded by Luna et al. (2013), is not fundamental. In contrast, it might be describing a specific time in the ever-changing state of X-ray-emitting symbiotic systems. The variations in the spectral properties within different epochs only mimic the extremely variable nature of these systems. The multi-epoch, multi-instrument observational data collected for T CrB allow us to peer into the dramatic changes that a symbiotic system experiences by evolving from a steady-state phase to an active one.

### v2

# The X-ray rise and fall of the Symbiotic Recurrent Nova System T CrB

@@ Abstract
We present the analysis of publicly available NuSTAR, Suzaku and XMM-Newton observations of the symbiotic recurrent nova T CrB covering the 2006.77–2022.66 yr period. The X-ray spectra are analysed by adopting a model that includes a reflection component produced by the presence of a disk that mimics the accretion disk and the immediate surrounding medium. Our best-fit model requires this disk to have a radius of 1 AU, effective thickness of 0.1 AU, averaged column density 1025 cm-2 and orientation of 50 with respect to the line of sight. This disk is about a factor of two larger than recent estimations for the accretion disk and its presence contributes significantly via reflection to the total X-ray flux detected from T CrB, which naturally produces the emission of the 6.4 keV Fe line. Our analysis suggests that the temperature of the boundary layer evolved from 14.8 keV in the steady-state phase (before 2016), to 2.8 keV in the 2017.24 epoch, to finally stabilise to about $\sim$8 keV in the subsequent epochs. These variations in the plasma temperature of the boundary layer are attributed to the evolution of the mass accretion rate ($\dot{M}_{\mathrm{acc}}$), which is estimated to have an averaged value of $\dot{M}_{\mathrm{acc}}$=2.6×$\times 10^{-8}$ M yr-1 for the current active phase. The presence of emission lines in the XMM-Newton RGS spectrum of 2017.24 prevents from adopting a black body emission model to fit the soft X-ray range. Instead, we use plasma emission models that suggest the presence of adiabatically-shocked gas produced by gas velocities of 110–200 km s-1, very likely tracing jet-like ejections similar to what is found in other symbiotic systems. The analysis of X-ray and optical data together show that T CrB has a similar evolution as black hole binaries, accreting neutron stars and AGN in the hardness-intensity diagram.
@@
pubyear: 2024pagerange: The X-ray rise and fall of the Symbiotic Recurrent Nova System T CrB8
@@ 1 Introduction
As a consequence of these works, we have started a series of papers to address the X-ray properties of the reflection component in highly variable symbiotic stars. We want to test further our ionised disk model to try to push forward our understanding of the accretion and reflection processes in symbiotic systems. Here we present the analysis of publicly available X-ray observations of T CrB. This is a recurrent symbiotic star in which the secondary is a M4III star (Mürset & Schmid, 1999). It has a period is 227.6 d (Kenyon & Garcia, 1986; Fekel et al., 2000) and its distance has been estimated to be $d$=896$\pm$22 pc using Gaia data (Bailer-Jones et al., 2021) [^1] Parallax distance measurements of symbiotic stars are affected by significant uncertainties given their extension, colours and orbital motion (see, e.g., Linford et al., 2019; Sion et al., 2019), however, with all this caveats, the one reported is the best available estimate of T CrB distance. Other distance estimates range up to 1 kpc (Hric et al., 1998), which might affect less than 25 percent the luminosity estimations.
@@
T CrB was first detected in X-rays by the Einstein observatory (Cordova et al., 1981) and has been routinely observed by more modern X-ray instruments. The first hard X-ray detection of T CrB during its quiescent state was reported by Tueller et al. (2005), but subsequent analysis of the Swift and Suzaku data suggested the presence of heavily-absorbed, high-temperature plasma, producing hard X-ray emission (Luna et al., 2008; Kennea et al., 2009). This source was then classified as a $\delta$-type symbiotic star, but by early 2014 it started brightening in the optical, entering a "super active" phase (Munari et al., 2016). In fact, the optical light curve exhibited a peak around 2016 April 4–5 (2016.35) and subsequent XMM-Newton observations obtained during 2017 showed that its spectrum developed a bright soft component, which changed T CrB classification into a $\beta/\delta$-type (Luna et al., 2018). Further analysis of 2017 and 2018 XMM-Newton observations in comparison with Suzaku and Swift data presented by (Zhekov & Tomov, 2019) showed that the flux of the soft component was declining while the opposite occurred to the hard component. Zhekov & Tomov (2019) estimated mass accretion rates ($\dot{M}_{\mathrm{acc}}$) for the different epochs and found that this was higher for the quiescent state than for the active phase (see fig. 6 of that publication).
@@
T CrB offers us a unique possibility to study the extreme changes of accretion states experienced by a symbiotic recurrent nova. In this paper, we present the analysis of multi-epoch (2006.77–2022.66), multi-instrument (XMM-Newton, Suzaku and NuSTAR) X-ray observations that cover the evolution from the steady-state phase up to the current active phase that, in combination with optical monitoring, can help us assess the role of the accretion process and the reflection physics. The analysis presented here is paving the road for the understanding of such an iconic recurrent symbiotic system on its way to producing a nova-like outburst in the near future.
@@ 2.1 XMM-Newton observations
Figure 2: $B$- (top panels) and $V$-magnitude (bottom panels) light curves of T CrB. Data from AAVSO, Schaefer (2010) and DASCH are illustrated with different colours. The data from Schaefer (2010) and DASCH have been displaced by 78.8 yr to fit the AAVSO light curve on year 2024.19. The vertical lines show the epoch of the X-ray observations analysed here. The grey- and yellow-shaded areas in the rightmost panels represents the possible next nova-like event predicted by Schaefer (2023) and Munari et al. (2016), respectively.
@@ 2.4 Complementary observations
We also retrieved the optical light curves of T CrB from the American Association of Variable Star Observers (AAVSO) [^6] https://www.aavso.org/ obtained with the optical $B$, $V$ and $I$ filters which are contemporaneous with the Observations obtained with the current generation of X-ray satellites. The data correspond to photometric measurements obtained between 2004 Mar 11 and 2024 June 11. The resultant light curves of the $B$ and $V$ filters are shown in Fig. 2 and a comparison between these two and the $I$ band photometry is presented in Appendix A. This figure also shows the photometric measurements obtained from the Digital Access to a Sky Century @ Harvard (DASCH) [^7] http://dasch.rc.fas.harvard.edu/lightcurve.php and values listed in Schaefer (2010). The latter data sets were shifted +78.8 yr to make them coincident with the optical dip observed on the year 2024.19 (marked with an arrow in the right panels of Fig. 2).
We note that this shift seems to suggests that T CrB will experience its next nova-like event on 2024.9 (about the end of October of the present year). Other predictions suggest this event to take place on 2024.4$\pm$0.4 (Schaefer, 2023) and 2025.5$\pm$1.3 (Munari et al., 2016). These two possible time ranges are illustrated in Fig. 2.
@@ 3 Data analysis and Results
It is interesting to note that the RGS spectrum obtained on 2017.24 shows the presence of emission lines (top right panel of Fig. 1). There seems to be emission from the C v, S xi, N vi, Ar xv and N vii at 35.0 Å, 31.5 Å, 29.1 Å, 27.05 Å  and 24.78 Å, respectively, which have been also detected in other X-ray-emitting symbiotic systems (see Toalá, Botello & Sabin, 2023; Toalá et al., 2023). Given that the RGS instruments are best suited to study the emission in the 0.33–2.5 keV energy range, the 2017.24 RGS spectrum is evidently tracing the contribution from the soft component. The estimated signal-to-noise of these emission lines range between 2.0 and 3.4 and their presence is also suggested by the medium-resolution EPIC MOS spectra (not shown here) but with evidently lower spectral quality. This suggests that the soft X-ray emission has a thermal origin (see also the cases of R Aqr and V694 Mon discussed; Sacchi et al., 2024; Lucy et al., 2020), instead of a black body as suggested in previous analysis of the XMM-Newton data (see Luna et al., 2018; Zhekov & Tomov, 2019).
@@
$<F_{V}>$ erg cm-2 s-1 6.7×$\times 10^{-10}$ 1.1×$\times 10^{-9}$ 1.4×$\times 10^{-9}$ 9.3×$\times 10^{-10}$ 9.6×$\times 10^{-10}$ 9.3×$\times 10^{-10}$ 9.1×$\times 10^{-10}$
$<F_{B}>$ erg cm-2 s-1 2.6×$\times 10^{-10}$ 6.0×$\times 10^{-10}$ 7.6×$\times 10^{-10}$ 4.6×$\times 10^{-10}$ 5.8×$\times 10^{-10}$ 5.0×$\times 10^{-10}$ 5.3×$\times 10^{-10}$
$<F_{I}>$ erg cm-2 s-1 4.0×$\times 10^{-9}$ 4.8×$\times 10^{-9}$ 5.2×$\times 10^{-9}$ 4.5×$\times 10^{-9}$ 4.2×$\times 10^{-9}$ 4.4×$\times 10^{-9}$ 4.0×$\times 10^{-9}$
$\dot{M}_{\mathrm{acc}}$ M yr-1 2.4×$\times 10^{-8}$ 4.1×$\times 10^{-8}$ 2.0×$\times 10^{-8}$ 2.4×$\times 10^{-8}$ 2.3×$\times 10^{-8}$ 2.4×$\times 10^{-8}$
@@ 3.1 Spectral analysis
The spectral analysis of the X-ray data of T CrB presented here was performed by means of the X-Ray Spectral Fitting Package (xspec; Arnaud, 1996). Following previous analysis of the X-ray data of $\delta$ and $\beta/\delta$ symbiotic systems, including T CrB (e.g., Zhekov & Tomov, 2019; Lucy et al., 2020), we decided to model the heavily-extinguished hot component by adopting an optically-thin emission model apec, included in xspec [^8] https://heasarc.gsfc.nasa.gov/xanadu/xspec/manual/node134.html. However, the presence of emission lines in the soft X-ray range unveiled by the 2017.24 data prevents us from modelling the soft range (0.3–2.0 keV) with a black body emission model. The soft X-ray emission was hence modelled with a second apec component. The extinction is taken into account by adopting photoelectric absorption using cross-sections of the phabs model, also part of xspec.
@@
Taking advantage of the analysis of the $\beta/\delta$ symbiotic system CH Cyg presented by our group (Toalá et al., 2023), we decided to use tailor-made reflection models in the spectral analysis of T CrB. We produced 200 disk models using the radiative-transfer code reflex (Paltani & Ricci, 2017) which simulates physical processes of X-ray photons propagating through matter around the central source, adopting Monte Carlo calculations to track individual photons. These models are characterised by an averaged column density $N_{\mathrm{H,ref}}$, a radius $R$, a thickness $h$ (such that $h=0.1R$) and a line of sight $\theta$. We sampled $N_{\mathrm{H,ref}}$=[1023,1024,1025,1026] cm-2 and $R$=[0.3, 0.5, 1.0, 10, 100] AU. For each disk model, we created 10 spectral energy distributions (SED) adopting angles between 0 and 90 (see illustration in Fig. 3). Finally, we converted these models to an additive one-parameter table (in fits format) associated with all the SEDs using the HEASoft task ftflx2tab [^9] https://heasarc.gsfc.nasa.gov/lheasoft/help/ftflx2tab.html.
It is important to mention here that these disk structures do not represent only the accretion disk around the WD component, they are meant to represent all the structures in the circumstellar medium around the symbiotic system (including the accretion disk). For example, numerical simulations presented in Liu et al. (2017) and de Val-Borro et al. (2017) show complex toroidal-like circumstellar material around the accreting WD. However, for simplicity we will refer to this model as reflection, because is the morphological feature that produces the reflection, particularly, its innermost region (see discussion section).
We adopted the following model configuration:
@@
We used the 2017.24 epoch to restrict $N_{\mathrm{H1}}$ and used this as a fixed value for the rest of the epochs. Table 2 shows that $N_{\mathrm{H1}}$ resulted in 2.3×$\times 10^{21}$ cm-2, which is very close to that estimated by Galactic H i column density towards T CrB (see section 3.2 of Zhekov & Tomov, 2019). The CF was left as a free parameter in the determination of the 2006.77+2015.81 epoch, which resulted in 0.998$\pm$0.001. This value was taken as a fixed parameter for the rest of the epochs. Fixed values are shown with boldface text in Table 2.
Figure 6: Temporal evolution of the X-ray flux and its components (top panel), mass accretion rate $\dot{M}_{\mathrm{acc}}$ (middle panel) and the ratio of the X-ray flux over that obtained from the optical $B$, $V$ and $I$ bands (bottom panels). The vertical dotted line shows the time of the maximum optical peak reported on 2016.35.
The temperature of the soft component ($kT_{1}$) is more or less consistent within different epochs, with an averaged value of 0.03 keV (=3.5×$\times 10^{5}$ K), however, its observed flux varies about two orders of magnitudes between epochs, being that of 2017.24 the brightest soft component. The heavily-extinguished apec ($kT_{2}$) and the reflection components jointly fit the unresolved Fe feature and, in the case of the 2006.77+2015.81 and 2020.77 epochs, the energy range above 10 keV detected by NuSTAR (see top right panel of Fig. 4 and top left panel of Fig. 5). The reflection model dominates for $E>$15 keV, leveraging the role of reflection in the hard energy range. The heavily-extinguished thermal component shows a temperature ranged from $kT_{2}$=2.8 keV (=3.2×$\times 10^{7}$ K) to values of $kT_{2}$=14.8 keV (=1.7×$\times 10^{8}$ K) with an hydrogen column density varying from $N_{\mathrm{H2}}$=25×$\times 10^{22}$ cm-2 to about 60×$\times 10^{22}$ cm-2. The top panel of Fig. 6 illustrates the variation of the total intrinsic flux $F_{\mathrm{X,TOT}}$ and the corresponding contributions from the three components ($F_{\mathrm{X,soft}}$, $F_{\mathrm{X,Hard}}$ and $F_{\mathrm{X,Ref}}$).
@@ 4 Discussion
As mentioned above, using the 2006.77+2015.81 spectrum it was possible to assign a $\delta$-type to T CrB (see table 1 in Merc et al., 2019), but any other latter medium-resolution X-ray spectrum can undoubtedly be used to classify this system as a $\beta/\delta$-type. This strongly suggest that the $\alpha$, $\beta$, $\delta$, and $\beta/\delta$ classification scheme, initially proposed by Mürset et al. (1997) and expanded by Luna et al. (2013), is not fundamental. In contrast, it might be describing a specific time in the ever-changing state of X-ray-emitting symbiotic systems. The variations in the spectral properties within different epochs only mimic the extremely variable nature of these systems. The multi-epoch, multi-instrument observational data collected for T CrB allow us to peer into the dramatic changes that a symbiotic system experiences by evolving from a steady-state phase to an active one.
4.1 The disk properties
Within the assumptions of the analysis presented here, the proposed disk model can reproduce satisfactorily the different epochs of observations. The best-fit is achieved with an inclination angle between the symmetry axis of the disk and the line of sight of $\theta=50^{\circ}$, which is very close to the value of 60$\pm$5 reported by Belczynski & Mikolajewska (1998).
Luna et al. (2018) estimated a radius for the accretion disk in the steady state of 0.75 R(=3.5×$\times 10^{-3}$ AU), however, a very recent optical monitoring of T CrB presented by Zamanov et al. (2024) concluded that its accretion disk has a size that ranges between 60 R(=0.28 AU) and 120 R(=0.56 AU), with an averaged value of 89$\pm$19 R(=0.42$\pm$0.09 AU). This value is about the same expected value as that of the Roche lobe of the WD component in T CrB. Our best fit model requires the disk structure to have a value twice as that estimated by Zamanov et al. (2024), $R$=1 AU. Which is not dramatically large if one takes into account the actual density distribution around the accreting WD in a symbiotic system as predicted by numerical simulations (see for example fig. 2 in Lee et al., 2022). Again, we remark that smaller disk structures resulted in unacceptable models because they do not produce the necessary reflection component to produce good fits.